HomeLife Life Insurance Company has two service departments (actuarial and premium rating) and two production departments (advertising and sales). The distribution of each service departments efforts (in percentages) to the other departments is shown in the following table: Premiumm ng Sales rom Actuarial- Premium 20% - 75% 10% 20 15% 60 The direct operating costs of the departments (including both variable and fixed costs) are: Actuarial Premium rating Advertising Sales S88,000 23,000 68,000 48,000 Required 1. Determine the totail cost allocated to the advertising and sales departments using the direct method Total Cost Allocated Advertising department Sales department 9102 20
media%2F503%2F503c5bc0-c108-4a78-9e3b-7c
Answer #1

1) Allocation of cost using Direct Method (Amounts in $)

Particulars Acturial Premium Advertising Sales
Direct Operating cost 88,000 23,000 68,000 48,000
Allocation of Acturial deptt cost in the ratio of 10:15 (88,000) - 35,200 52,800
Allocation of Premium deptt cost in the ratio of 20:60 - (23,000) 5,750 17,250
Total cost allocated 0 0 108,950 118,050

2) Allocation of cost using Step Method (Amounts in $)

Particulars Acturial Premium Advertising Sales
Direct Operating cost 88,000 23,000 68,000 48,000
Allocation of Acturial deptt cost in the ratio of 75:10:15 (88,000) 66,000 8,800 13,200
Allocation of Premium deptt cost in the ratio of 20:60 - (89,000) 22,250 66,750
Total cost allocated 0 0 99,050 127,950

3) Let the Acturial department denoted by A and Premium department by P.

A = $88,000+(0.20*P) - Eqn 1

P = $23,000+(0.75*A) - Eqn 2

Now put the value of P in eqn 1 which is shown as follows:-

A = 88,000+[0.20*(23,000+0.75A)]

A = 88,000+4,600+0.15A

A - 0.15A = 92,600

A = 92,600/0.85 = 108,941

Therefore total cost allocated to Acturial department is $108,941

Total cost allocated to Premium department = $23,000+(0.75*A)

= $23,000+($108,941*0.75)

= $23,000+$81,705 = 104,705

Allocation of cost using Reciprocal Method (Amounts in $)

Particulars Acturial Premium Advertising Sales
Direct Operating cost 88,000 23,000 68,000 48,000
Allocation of Acturial deptt cost in the ratio of 75:10:15 (108,941) 81,705 10,895 16,341
Allocation of Premium deptt cost in the ratio of 20:20:60 20,941 (104,705) 20,941 62,823
Total cost allocated 0 0 99,836 127,164